OpenStudy (anonymous):

how do you solve this : m^2 -2m+1= 9

Nnesha (nnesha):

set it equal to 0 so subtract 9 both sides

Nnesha (nnesha):

\[\huge\rm m^2 -2m +1 \color{Red}{-9} =0\] 1-9 = -8 \[\huge\rm m^2 -2m -8 =0\] now find two number if you multiply them you should get product of AC(a is leading coefficient and C is constant term of the equation ) and if you add or subtract them you should get middle term a b c are \[\large\rm \color{ReD}{A}x^2 + \color{red}{B}x+\color{red}{C}=0\]

Nnesha (nnesha):

\[\large \rm \color{green}{1}m^2 \color{green}{-2}m \color{green}{-8} =0\]\[\large\rm \color{ReD}{A}x^2 + \color{red}{B}x+\color{red}{C}=0\]

OpenStudy (anonymous):

m = -2 or 4

Nnesha (nnesha):

@RKSSA sorry but your answer is wrong ._.